# SPICE: A Synergistic, Precise, Iterative, and Customizable Image Editing Workflow

### v1

@@ Appendix B Human Evaluation Details
An example of an image pair the annotator sees is shown in Figure 15. To ensure fairness, we send the same instructions to the annotators. To ensure minimal bias, we do not explain the relative strengths and weaknesses of each method. We do not further communicate with annotators or provide clarifications in written or spoken form, until the annotators finishes the task. The only message we send to annotators is shown below (from “Hi” to “winner”).
@@
Figure 15: An example of an image pair that the annotator sees. The annotator does not know which option comes from which model. While they can see the name of the baseline method they are evaluating, none of the annotators know about the baseline methods.
Appendix C Challenging Examples from More Benchmarks
We further compare the performance of our workflow against baseline methods on selected challenging examples. A challenging example is an example where all baseline methods fail. We select challenging examples from the Emu Edit and MagicBrush test sets. For the Emu Edit test set, the baseline methods are Emu Edit, IP2P, MagicBrush, and UltraEdit (Figure 16 and Figure 17). For the MagicBrush test set, the baseline methods are Reference (the best DALL·E 2 generations), IP2P, MagicBrush, and UltraEdit (Figure 18 and Figure 19). After we exclude the style change task, both test sets cover 7 different editing categories, as reported in the original papers. We select 2 challenging examples from each category, totaling 14 for each test set. Our workflow succeeds on all challenging examples. Our workflow also outperforms SeedEdit from Doubao AI on these examples. The authors translate the original prompts into Chinese when using SeedEdit.
Original Emu Edit IP2P MagicBrush UltraEdit Doubao Ours
@@ Appendix C Challenging Examples from More Benchmarks
Figure 16: Our workflow performs the best on the first half of 14 challenging examples from Emu Edit. Doubao AI refuses to generate an image for the washing machine example, showing an error message of “Sorry, I cannot generate the image you requested”.
Original Emu Edit IP2P MagicBrush UltraEdit Doubao Ours
@@
Original Reference IP2P MagicBrush UltraEdit Doubao Ours
@@
Original Reference IP2P MagicBrush UltraEdit Doubao Ours
@@
Appendix D Ablation Studies
We qualitatively demonstrate the effect of removing each component from our workflow (LABEL:fig:ablation). Then, we demonstrate that our workflow mitigates many issues of the specialized inpainting model FLUX.1 [dev] Fill (LABEL:fig:flux-fill). Together, the analysis in this section shows that our workflow improves upon the backbone image generation model.
D.1 Necessity of Components
Removing any component from our workflow leads to the following negative consequences (LABEL:fig:ablation).
@@ D.1 Necessity of Components
D.2 Flux.1 [dev] Fill Failures
Our workflow mitigates many issues of the Flux.1 [dev] Fill, a specialized inpainting model trained on the same backbone of Flux.1 [dev]. We discuss the issues below and demonstrate them in LABEL:fig:flux-fill.
@@
Flux.1 [dev] Fill sometimes results in an obvious color drift outside the mask (saturation of the grass color drops), a weakness acknowledged by the official developers [^18] https://huggingface.co/black-forest-labs/FLUX.1-Fill-dev#limitations. This is because unlike our workflow, Flux.1 [dev] Fill is not strictly a local editing model. By design, our method strictly preserves pixel values outside the mask.
@@
Appendix E Measuring Properties of Generated Objects
@@
Appendix F Changing the Random Seed
@@
With baseline methods, a user can change the random seed to get multiple results and select the best one. However, even with more compute, baseline models will consistently fail on the same challenging example (LABEL:fig:change-seed). In contrast, with our workflow, a user can better utilize the increased compute if the user performs editing step by step.
Appendix G Iterative Construction
@@
Figure 33 shows all 40 steps we use to iteratively construct LABEL:subfig:iterative-heart. Some steps involve subtle but important edits that are best visualized at a high resolution. We will release a video to better illustrate the editing operation done at each step.
Appendix H Results of Different Art Styles
@@
Figure 34 shows two additional results. One is generated with BoleroMix (Pony) v1.41 as the backbone, and the other is generated with the same Flux.1 [dev] as the backbone. These additional results illustrate the flexibility of our workflow in editing and iteratively generating images of different art styles. Note that these are fan-made images, where details may differ from those of the original characters.
@@
BoleroMix (Pony) v1.41 (Characters from Touhou Project) Flux.1 [dev] (Characters from Hades 2)
@@
Figure 34: Our workflow excels at editing and iteratively generating images of different art styles. Both images are generated using around 100 to 200 editing steps, within 4 hours in one sitting. Character LoRAs are not used, so many editing steps are spent on correcting details of the character’s clothes. Note that due to the complex composition and high resolution (2048$\times$2048), it is extremely hard, if not impossible, to generate these images with a single text-to-image step.
Appendix I Hyperparameter Recommendations
To generate LABEL:subfig:customizable, the recommended hyperparameter ranges are shown in Table 2. Note that several consecutive editing steps are necessary in a certain editing region to achieve the shown effect.
@@ Appendix I Hyperparameter Recommendations
Appendix J Comparison with MagicQuill
MagicQuill (Liu et al., 2024b) is a method very similar to ours, because it also accepts edge and color information from the user as its input. However, the following differences in design choices lead to a better performance of our method.
@@
As our workflow uses different models for different stages, it can be adapted even when the ControlNets are not released as a module on the base model but as a full model (for example, the official Flux.1 [dev] Canny model by Black Forest Labs). Hence, the higher flexibility allows us to adapt our workflow to the Flux.1 [dev] base model within one week of the release of Flux.1 [dev] Canny in November 2024, whereas MagicQuill still supports neither SDXL or FLUX models by April 2025202020https://github.com/ant-research/MagicQuill/issues/89.
@@
Our workflow does not downsample the color information. Hence, high frequency details are preserved from user input. In contrast, MagicQuill downsamples user color input to a 32$\times$32 resolution during pre-processing, an information bottleneck that limits the ability of users to provide detailed color hints.
@@
MagicQuill assumes that the user only wants to edit the region around locations where the color patch is provided. The mask is uniformly expanded in all directions from the color patch. However, consider the example of adding a backpack on a bench (Figure 35). In this case, the mask should be expanded more in the horizontal direction than the vertical direction, because the mask should not fully cover the backrest of the bench. To address this difficulty, the disentangled design in our workflow returns the freedom to the user.
@@
By changing parameters and resolutions, our workflow smoothly transitions from an image editor to a detail enhancer, or from an inpainter to an outpainter. Examples of both transitions can be found in Figure 33. Neither detail enhancing nor outpainting is supported by MagicQuill.
Due to the above advantages, our method outperforms MagicQuill on a challenging example (Figure 35), with the same Realistic Vision V6.0 B1 (based on SD 1.5) model [^21] https://civitai.com/models/4201?modelVersionId=245598 as the backbone. To ensure fair comparison, we take the first-shot, single-step result from our method, but try our best to tune the hyperparameters and to draw accurate shapes for MagicQuill. After over 20 attempts, we are still unable to correctly add the backpack onto the bench using MagicQuill.
@@
Our Hinted Image Our Blurred Mask Our Result (SD 1.5)
Refer to caption Refer to caption Refer to caption
MagicQuill Gemini 2.0 Flash GPT-4o
Refer to caption Refer to caption Refer to caption
Figure 35: Using Realistic Vision V6.0 B1, an SD-1.5-based model released in December 2023, as the backbone, our workflow outperforms both MagicQuill (with the same backbone) and latest commercial VLMs (released in March 2025). For our workflow and MagicQuill, we used the same description prompt “black backpack on a bench”. For Gemini 2.0 Flash and GPT-4o, we used the same editing prompt “add a black backpack onto this bench”. Only our method results in almost perfect spatial relationships, with only a minor error at the bottom of the backpack. MagicQuill is unable to generate the lower part of the backpack, which is supposed to be visible through slits on the backrest. Gemini 2.0 Flash and GPT-4o both generate a floating backpack in front of the backrest. The original image and our better result of using Flux.1 [dev] as the backbone are shown in Figure 3.
Appendix K Limitations
@@
If the base model has some strong local minima, our workflow is unable to help the user generate content out of those local minima. Two examples from the base model (Flux.1 [dev] with a LoRA) we use are (1) a fixed hairstyle regardless of the edge hints and (2) two patellas on one knee regardless of the color hints. To mitigate these artifacts, a convenient and successful workaround is to increase the Canny model steps (correspondingly reducing the base model steps), since the Canny model and the base model seldom suffer from the same local minima. The resulting image after successful edits is shown in the right panel of Figure 34.
@@
Appendix L Workflow Implementation
@@
Other than Flux.1 [dev], we thoroughly test our workflow with BoleroMix (Pony) v1.41222222https://civitai.com/models/448716?modelVersionId=629179, a Japanese animation-style base model derived from SDXL (Podell et al., 2024). For this one and other SDXL-derived models, the user can implement the workflow by simply enabling relevant add-ons in Stable Diffusion Web UI Automatic1111. We also test our workflow in other models at a smaller scale, where all steps in our workflow similarly work.
We have tested our workflow on one NVIDIA RTX A6000, NVIDIA A100 (80 GB), or NVIDIA GeForce RTX 4090. On A6000, each editing step using Flux.1 [dev] (1024$\times$1024 resolution and 30 diffusion steps) only takes around 30 seconds. The price of a server with a single A6000 can be as low as 0.5\$ per hour [^23] https://cloud.vast.ai/ (January 2025), equivalent to 0.004\$ per editing step. This is one tenth of the cost of DALL·E 3, which is 0.04\$ per image [^24] https://openai.com/api/pricing/. Due to the low success rate of DALL·E 3 on hard editing tasks (Section 3.4), the cost gap is only larger in real scenarios.

### v2

@@ Appendix D Human Evaluation Details
An example of an image pair the annotator sees is shown in Figure˜11. To ensure fairness, we send the same instructions to the annotators. To ensure minimal bias, we do not explain the relative strengths and weaknesses of each method. We do not further communicate with annotators or provide clarifications in written or spoken form, until the annotators finishes the task. The only message we send to annotators is shown below (from “Hi” to “winner”).
@@
Figure 11: An example of an image pair that the annotator sees. The annotator does not know which option comes from which model. While they can see the name of the baseline method they are evaluating, none of the annotators know about the baseline methods before performing annotation.
Appendix E Challenging Examples from More Benchmarks
We further compare the performance of our workflow against baseline methods on selected challenging examples. A challenging example is an example where all baseline methods fail. We select challenging examples from the Emu Edit and MagicBrush test sets. For the Emu Edit test set, the baseline methods are Emu Edit, IP2P, MagicBrush, UltraEdit, Doubao SeedEdit, Gemini 2.0 Flash, and GPT-4o (Figure˜12 and Figure˜13). For the MagicBrush test set, the baseline methods are Reference (the best DALL·E 2 generations), IP2P, MagicBrush, UltraEdit, Doubao SeedEdit, Gemini 2.0 Flash, and GPT-4o (Figure˜14 and Figure˜15). After we exclude the style change task, both test sets cover 7 different editing categories, as reported in the original papers. We select 2 challenging examples from each category, totaling 14 for each test set. We translate the original prompts into Chinese when using Doubao SeedEdit.
Our workflow qualitatively outperforms strong baselines including Doubao SeedEdit, Gemini 2.0 Flash, and GPT-4o. Notably, our workflow strictly preserves details outside the edited region. In contrast, fine-grained texture of the image corrupt severely after a single step of editing by either Gemini 2.0 Flash or GPT-4o, which is an innate limitation of mask-free methods. Moreover, GPT-4o stretches a square input image into a rectangular image without being prompted, leading to inconsistent resolutions.
Original Emu Edit IP2P MB UE Doubao Gemini GPT-4o Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@ Appendix E Challenging Examples from More Benchmarks
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 12: Our workflow performs the best on the first half of 14 challenging examples from Emu Edit. Both Doubao AI and GPT-4o refuses to generate an image for the washing machine example, showing an error message instead.
Original Emu Edit IP2P MB UE Doubao Gemini GPT-4o Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Original Reference IP2P MB UE Doubao Gemini GPT-4o Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Original Reference IP2P MB UE Doubao Gemini GPT-4o Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Appendix F Ablation Studies
We qualitatively demonstrate the effect of removing each component from our workflow (Figure˜16). Then, we demonstrate that our workflow mitigates many issues of the specialized inpainting model FLUX.1 [dev] Fill (Figure˜17). Together, the analysis in this section shows that our workflow improves upon the backbone image generation model.
F.1 Necessity of Components
Removing any component from our workflow leads to the following negative consequences (Figure˜16).
@@ F.1 Necessity of Components
F.2 Flux.1 [dev] Fill Failures
Our workflow mitigates many issues of the Flux.1 [dev] Fill, a specialized inpainting model trained on the same backbone of Flux.1 [dev]. We discuss the issues below and demonstrate them in Figure˜17.
@@
Flux.1 [dev] Fill sometimes results in an obvious color drift outside the mask (saturation of the grass color drops), a weakness acknowledged by the official developers. [^21] https://huggingface.co/black-forest-labs/FLUX.1-Fill-dev#limitations This is because unlike our workflow, Flux.1 [dev] Fill is not strictly a local editing model. By design, our method strictly preserves pixel values outside the mask.
@@
Original Hinted No Context No Blur No Hints No Canny Full
Figure 16: Ablation studies show that each component of our workflow is necessary. Removing any one leads to suboptimal quality.
Original Flux.1 [dev] Fill Ours
Refer to caption Refer to caption Refer to caption
(a) Global Editing Failures: A person sits in a bedroom.
Original Flux.1 [dev] Fill Ours
Refer to caption Refer to caption Refer to caption
(b) Color Drifts: Two horses grazing on a grassy field under a clear sky.
Original Flux.1 [dev] Fill Ours
Refer to caption Refer to caption Refer to caption
(c) Prompt Bleeding: A flamingo standing on a camel walking on a desert.
Figure 17: Our workflow mitigates issues of the strong Flux.1 [dev] Fill model. While Flux.1 [dev] Fill is a specialized inpainting checkpoint, it suffers intrinsic limitations including failures in global editing, color drifts, and prompt bleeding. In contrast, our training-free workflow is free from these issues. In the captions, we show the description prompt for each editing task. Note that we do not use the editing prompt (“add a flamingo on top of the camel”), because these two methods are both designed to accept description prompts instead of editing prompts.
Appendix G Simple Inputs
While our workflow accepts extra inputs from the user, generating the extra inputs poses minimal burden. In Figure˜18, we show that simple color hints (Hinted) and simple masks (Mask) can produce high-quality edits.
Original Hinted Silhouette Mask Result
Appendix H Measuring Properties of Generated Objects
@@
Appendix I Changing the Random Seed
@@
With baseline methods, a user can change the random seed to get multiple results and select the best one. However, even with more compute, baseline models will consistently fail on the same challenging example (Figure˜19). In contrast, with our workflow, a user can better utilize the increased compute if the user performs editing step by step.
Seed 0 Seed 1 Seed 2 Seed 3 Seed 4 Seed 5 Seed 6 Seed 7 Seed 8 Seed 9
(a) IP2P
(b) MagicBrush
(c) UltraEdit
Figure 19: For baseline methods, changing the random seed does not help. The random seeds are shown on the top. With 10 times the original compute, baseline methods still fail on the challenging editing task from EditEval of adding a bag to the bench. Our workflow succeeds in the first edit and will succeed in fewer than 10 edits should the first edit fail. The original image and our result are in Figure˜5.
Appendix J Iterative Construction
@@
Figure˜24 shows all 40 steps we use to iteratively construct Figure˜9(a). Some steps involve subtle but important edits that are best visualized at a high resolution. We will release a video to better illustrate the editing operation done at each step.
We further evaluate iterative editing on recent (October 2025) state-of-the-art models and qualitatively demonstrate how image quality worsens as edits accumulate. The three models we evaluate are FLUX.1 Kontext [pro], Sora (GPT), and Gemini 2.5 Flash Image (Nano Banana). Access to FLUX.1 Kontext [Pro] and Sora was provided by their respective official websites, and access to Gemini 2.5 Flash Image (Nano Banana) was provided via Google AI Studio. [^23] https://flux1.ai and https://sora.chatgpt.com For each model, we apply a sequence of 12 iterative edits to the same base image of a man sitting on a chair, wearing a yellow jacket. [^24] https://www.pexels.com/photo/man-in-yellow-long-sleeve-shirt-sitting-on-black-chair-smiling-7562191/ The editing prompt at each step was: “Change the man’s jacket color to x” where x cycles twice through the colors red, orange, yellow, green, blue, and purple, for a total of 12 editing steps. All models failed in this iterative editing task, with image quality dropping dramatically after 12 steps. Below, we further explain and visualize the failure patterns of each model.
FLUX.1 Kontext [pro].
This model is a diffusion-based image editor with strong local inpainting and instruction-following capabilities [2]. However, repeated edits introduce obvious artifacts, as shown in Figure˜20.
Step 1 Step 2 Step 3 Step 4 Step 5 Step 6
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Step 7 Step 8 Step 9 Step 10 Step 11 Step 12
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 20: Noise accumulates in the iterative editing results by FLUX.1 Kontext [pro]. Obvious artifacts appear at around 7 steps.
Sora.
Sora preserves global structure well in early steps but changes the aspect ratio from the first step. Moreover, degradation appears as gradual over-smoothing and inconsistency in fine detail, as shown in Figure˜21. It also struggles with preserving facial structure as noted in practitioner reports. [^25] https://davidharris.uk/2025/04/11/kling-vs-runway-vs-sora-vs-lumalabs-a-deep-dive-into-ai-video-generation-platforms/
Step 1 Step 2 Step 3 Step 4 Step 5 Step 6
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Step 7 Step 8 Step 9 Step 10 Step 11 Step 12
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 21: Sora does not preserve details over 12 iterative editing steps. The background color changes from gray to light blue. The facial structure of the man also changes.
Gemini 2.5 Flash Image.
During iterative editing, Gemini 2.5 Flash Image preserves global structure reliably, but texture and details still degrade, as shown in both Figure˜22 and Figure˜23.
Step 1 Step 2 Step 3 Step 4 Step 5 Step 6
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Step 7 Step 8 Step 9 Step 10 Step 11 Step 12
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 22: Gemini 2.5 Flash Image fails to maintain character consistency over 12 editing steps. The background, skin tone, and details on the jacket all change.
Figure 23: For clarity, we present a full-page comparison between the initial image and the final output from Gemini 2.5 Flash Image. The skin tone shifts noticeably, and the background becomes darker and more textured. The buttons on the man’s jacket disappear in the edited version, and fine details (eyes, teeth, and fingernails) morph and differ from the original. Additional indentations and fine lines also appear around the mouth.
Appendix K Results of Different Art Styles
@@
Figure˜25 shows two additional results. One is generated with BoleroMix (Pony) v1.41 as the backbone, and the other is generated with the same Flux.1 [dev] as the backbone. These additional results illustrate the flexibility of our workflow in editing and iteratively generating images of different art styles. Note that these are fan-made images, where details may differ from those of the original characters.
@@
BoleroMix (Pony) v1.41 Flux.1 [dev]
@@
Figure 25: Our workflow excels at editing and iteratively generating images of different art styles. Both images are generated using around 100 to 200 editing steps, within 4 hours in one sitting. The characters are from Touhou Project and Hades 2. Character LoRAs are not used, so many editing steps are spent on correcting details of the character’s clothes. Note that due to the complex composition and high resolution ($\times2048$), it is extremely hard, if not impossible, to generate these images with a single text-to-image step.
Appendix L Hyperparameter Recommendations
We first demonstrate the effect of each hyperparameter (Figure˜26). Then, we recommend several combination of hyperparameters can help the user reliably generate an image that DALL·E 3 and GPT-4o cannot (Figure˜10).
Context Size.
During inpainting, a diffusion model generates content consistent with what is already in the context. In Figure˜26(a), the user requires another yellow vegetable to be added to the center of the image. The prompt is “yellow vegetables on an orange background”. The color hint is the yellow circle in the center. The model’s interpretation of the color hint depends on the selected context. When a small context includes none of the other vegetables, the model generates multiple vegetables from the single yellow circle. When a medium context includes corners of the other vegetables, the model generates one vegetable. When a large context includes all other vegetables, the model fails to generate a new vegetable. When the context includes one vegetable either on the top left or the top right corner, the model generates a vegetable with the same direction of shade. While we do not theoretically explain the failure when using large context sizes, we observe it frequently in other editing tasks. We also observe the failure to add objects using the “whole image” mode (Section˜2.1), an observation that motivated our design of context dots.
Denoising Strength.
The model balances hinted colors and realistic shadows at a medium denoising strength. In Figure˜26(b), the user requires a Rubik’s cube with certain face colors to be added to the image. Because the colors cannot be succinctly described by words, the user inserts a reference image as the hint. With a low denoising strength, the colors are preserved, but the shadows are not generated. We observe the opposite with a high denoising strength.
Canny Model Steps.
The model balances realistic details and hinted edges at a medium number of Canny model steps. In Figure˜26(c), the user requires a crescent to be added to the sky, but the prompt is simply “a moon”. With few Canny model steps, the moon looks realistic, but its shape is wrong. We observe the opposite with many Canny model steps.
Hinted Small Medium Large Top Left Top Right
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) Context Size
Hinted 0.1 0.3 0.5 0.7 0.9
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(b) Denoising Strength
Hinted 5 10 15 20 25
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(c) Canny Model Steps
Figure 26: Users can achieve desired effects by customizing the value of 3 hyperparameters. (a) To improve the consistency of edited region with a certain part of the image, the user can cover this part in the context. (b) Lower denoising strength preserves the color, whereas higher denoising strength inserts realistic shadows. (c) Fewer Canny model steps produce high-frequency details, whereas more steps allows the generated image to faithfully follow the hinted shape. Some examples look sub-optimal, because they show the results of only varying one hyperparameter. In practice, users can adjust all three parameters together for better edits.
To generate Figure˜10(c), the recommended hyperparameter ranges are shown in Table˜3. Note that several consecutive editing steps are necessary in a certain editing region to achieve the shown effect.
@@
Appendix M Comparison with MagicQuill
MagicQuill [14] is a method very similar to ours, because it also accepts edge and color information from the user as its input. However, the following differences in design choices lead to a better performance of our method.
@@
As our workflow uses different models for different stages, it can be adapted even when the ControlNets are not released as a module on the base model but as a full model (for example, the official Flux.1 [dev] Canny model by Black Forest Labs). Hence, the higher flexibility allows us to adapt our workflow to the Flux.1 [dev] base model within one week of the release of Flux.1 [dev] Canny in November 2024, whereas MagicQuill still supports neither SDXL or FLUX models by April 2025.262626https://github.com/ant-research/MagicQuill/issues/89
@@
Our workflow does not downsample the color information. Hence, high frequency details are preserved from user input. In contrast, MagicQuill downsamples user color input to a $\times32$ resolution during pre-processing, an information bottleneck that limits the ability of users to provide detailed color hints.
@@
MagicQuill assumes that the user only wants to edit the region around locations where the color patch is provided. The mask is uniformly expanded in all directions from the color patch. However, consider the example of adding a backpack on a bench (Figure˜27). In this case, the mask should be expanded more in the horizontal direction than the vertical direction, because the mask should not fully cover the backrest of the bench. To address this difficulty, the disentangled design in our workflow returns the freedom to the user.
Our Hinted Image Our Blurred Mask Our Result (SD 1.5)
Refer to caption Refer to caption Refer to caption
MagicQuill Gemini 2.0 Flash GPT-4o
Refer to caption Refer to caption Refer to caption
Figure 27: Using Realistic Vision V6.0 B1, an SD-1.5-based model released in December 2023, as the backbone, our workflow outperforms both MagicQuill (with the same backbone) and latest commercial VLMs (released in March 2025). For our workflow and MagicQuill, we used the same description prompt “black backpack on a bench”. For Gemini 2.0 Flash and GPT-4o, we used the same editing prompt “add a black backpack onto this bench”. Only our method results in almost perfect spatial relationships, with only a minor error at the bottom of the backpack. MagicQuill is unable to generate the lower part of the backpack, which is supposed to be visible through slits on the backrest. Gemini 2.0 Flash and GPT-4o both generate a floating backpack in front of the backrest. The original image and our better result of using Flux.1 [dev] as the backbone are shown in Figure˜5.
@@
By changing parameters and resolutions, our workflow smoothly transitions from an image editor to a detail enhancer, or from an inpainter to an outpainter. Examples of both transitions can be found in Figure˜24. Neither detail enhancing nor outpainting is supported by MagicQuill.
Due to the above advantages, our method outperforms MagicQuill on a challenging example (Figure˜27), with the same Realistic Vision V6.0 B1 (based on SD 1.5) model [^27] https://civitai.com/models/4201?modelVersionId=245598 as the backbone. To ensure fair comparison, we take the first-shot, single-step result from our method, but try our best to tune the hyperparameters and to draw accurate shapes for MagicQuill. After over 20 attempts, we are still unable to correctly add the backpack onto the bench using MagicQuill.
@@
Appendix N Limitations
@@
If the base model has some strong local minima, our workflow is unable to help the user generate content out of those local minima. Two examples from the base model (Flux.1 [dev] with a LoRA) we use are (1) a fixed hairstyle regardless of the edge hints and (2) two patellas on one knee regardless of the color hints. To mitigate these artifacts, a convenient and successful workaround is to increase the Canny model steps (correspondingly reducing the base model steps), since the Canny model and the base model seldom suffer from the same local minima. The resulting image after successful edits is shown in the right panel of Figure˜25.
@@
Appendix O Workflow Implementation
@@
Other than Flux.1 [dev], we thoroughly test our workflow with BoleroMix (Pony) v1.41, [^28] https://civitai.com/models/448716?modelVersionId=629179 a Japanese animation-style base model derived from SDXL [18]. For this one and other SDXL-derived models, the user can implement the workflow by simply enabling relevant add-ons in Stable Diffusion Web UI Automatic1111. We also test our workflow in other models at a smaller scale, where all steps in our workflow similarly work.
We have tested our workflow on one NVIDIA RTX A6000, NVIDIA A100 (80 GB), or NVIDIA GeForce RTX 4090. On A6000, each editing step using Flux.1 [dev] ($\times1024$ resolution and 30 diffusion steps) only takes around 30 seconds. The price of a server with a single A6000 can be as low as 0.5\$ per hour, [^29] https://cloud.vast.ai/ (January 2025) equivalent to 0.004\$ per editing step. This is one tenth of the cost of DALL·E 3, which is 0.04\$ per image. [^30] https://openai.com/api/pricing/ Due to the low success rate of DALL·E 3 on hard editing tasks (Section˜B.4), the cost gap is only larger in real scenarios.